# ChEX: Interactive Localization and Region Description in Chest X-rays

### v1

@@ Abstract
Report generation models offer fine-grained textual interpretations of medical images like chest X-rays, yet they often lack interactivity (i.e. the ability to steer the generation process through user queries) and localized interpretability (i.e. visually grounding their predictions), which we deem essential for future adoption in clinical practice. While there have been efforts to tackle these issues, they are either limited in their interactivity by not supporting textual queries or fail to also offer localized interpretability. Therefore, we propose a novel multitask architecture and training paradigm integrating textual prompts and bounding boxes for diverse aspects like anatomical regions and pathologies. We call this approach the Chest X-Ray Explainer (ChEX). Evaluations across a heterogeneous set of 9 chest X-ray tasks, including localized image interpretation and report generation, showcase its competitiveness with SOTA models while additional analysis demonstrates ChEX’s interactive capabilities.
@@
Vision-Language Modeling Localized Report Generation1 Introduction
@@ 2.0.2 Prompt-based Localization
DETR [6] pioneered the use of token vectors for object detection, sparking a series of subsequent models following this approach [91, 90, 41, 47, 86]. However, these models typically employ tokens that are not input-dependent, often relying on learned or position-based tokens [41, 86, 47]. Visual grounding models [9, 8, 11, 37] predict bounding boxes for given textual phrases, relating them to image regions, and have been applied in medical imaging [7, 27]. GLIP [35, 87] unifies object detection and visual grounding using textual prompts, with applications emerging in the medical domain [17, 81]. Open-vocabulary object detection extends object detection models to unseen classes using textual prompts [80, 16, 85, 42, 62, 46]. Similarly, prompt-based segmentation techniques have been explored [33] with applications in the medical domain [25, 45].
@@ 3 ChEX: A Localized and Interactive Chest X-ray Description Model
Our model ChEX supports a wide range of tasks, spanning from localization to text generation, while considering different aspects like anatomical regions and pathologies. To facilitate efficient training and zero-shot inference across all tasks, ChEX employs a simple yet versatile architecture, outlined in Figure 2. First, an image encoder is used to extract patch tokens of the given chest X-ray, while each textual prompt is encoded using the prompt encoder, a frozen text encoder. Next, the prompt detector, a DETR[6]-style object detector, localizes the prompts within the image, predicting a set of bounding boxes for each prompt along with a single Region-of-interest (ROI) token vector per prompt. Finally, the sentence generator is conditioned on each ROI token independently as well as on all patches to predict a concise description of each ROI. Further details are given in the following Secs. 3.1, 3.2 and 3.3. For reproducibility, we provide comprehensive implementation details in Appendix 0.D.
@@ 3.2 Multitask Training
We train ChEX in a multitask setting where each sample may provide targets for different aspects of the image. To enable training in such a setting, we use three different types of queries depending on the available targets per sample: (i) pathology tokens, (ii) anatomy tokens, and (iii) sentence tokens.
Pathology tokens are employed when pathology class labels or bounding boxes are available with a sample. We encode the textual prompt of each pathology, which then serves as a prompt token for the prompt detector. The generated ROI tokens are trained via a (supervised) contrastive loss, incorporating both positive and negative prompts for each pathology. In positive pairs, the ROI token for a pathology is paired with its positive or negative pathology prompt, depending on the presence or absence of this pathology. Conversely, negative pairs include the opposing prompt alongside all positive and negative prompts of all other pathologies. If target bounding boxes are available, we use a modified version of the DETR loss [6]. We iteratively apply Hungarian matching per pathology until all predicted boxes are matched and then apply the L1 and gIoU losses from the original DETR loss. Box scores are trained by applying the focal loss [39] with positive targets for boxes matched in the first iteration.
Anatomy tokens are employed when anatomical region targets are available, including bounding boxes, associated pathology labels, and report sentences per region. We use the anatomical region names as textual prompts, training their bounding boxes using L1 and gIoU losses, while box scores remain untrained. If multiple target boxes are available per region, we apply Hungarian matching. For regions with a single bounding box, we match all its predicted boxes with the target box. Similar to the approach adopted for pathological tokens, we apply a supervised contrastive loss using positive and negative pathology prompts. For positive pairs, we align the ROI token of a region with the set of all positive pathology prompts of pathologies present in this region, along with negative prompts of pathologies not present. All other pathology prompts are used as negatives. Additionally, we train the sentence generator conditioned on anatomy features (i.e. ROI tokens) using the target sentences assigned to each anatomical region. We apply the standard autoregressive language modeling loss considering only regions where target sentences are available. If multiple target sentences are available, we concatenate them. For each sample, we randomly decide whether the anatomy features (i.e. ROI tokens) are computed using textual anatomy prompts or queries based on target bounding boxes of anatomical regions.
Sentence tokens are employed when report sentences are available. In such cases, the encoded sentences serve as prompt tokens. Since target bounding boxes and pathology labels are absent for sentences, we rely solely on the sentences themselves. We use a contrastive loss, where the extracted ROI token for each sentence and the encoded sentence itself form positive pairs, while other encoded sentences from the same and different samples serve as negatives. Additionally, we apply the standard CLIP loss [58] on average-pooled patch and sentence features. Furthermore, we train the sentence generator to reconstruct the original sentence based on its ROI token.
@@ 4.2 Benchmark Tasks
We evaluate the zero-shot performance of our model ChEX across the 9 chest X-ray tasks shown in Tab. 1. For details on the evaluation and dataset preparation, we refer to Appendix 0.E. While no task-specific fine-tuning is applied, post-processing (e.g. box suppression and scaling of boxes) for ChEX and all baselines is adjusted to consider differences in annotation practices of datasets.
@@ 5 Results
We compare ChEX with SOTA models (Sec. 5.1). Additionally, we study how ChEX reacts to (interactive) user queries (Secs. 5.2 and 5.3) and show its interpretable and customizable report generation capabilities (Sec. 5.4). ChEX is competitive with SOTA models while providing a high degree of interpretability and interactivity, therefore offering a promising path towards clinical application.
@@ 5.1 Comparison with SOTA
Fig. 3 and Tab. 2 provide an overview of the performance of our model ChEX compared to the best baselines, including specialized SOTA and common multitask models. On 8 of the 9 tasks, our model ChEX is competitive (within 1-std) or better than the best baseline on at least one metric. Only on pathology detection (OD) on VinDR-CXR is it outperformed by a specialized supervised object detection model. Note that none of the baselines is capable of performing all the tasks, as most of them are either focused solely on localization or generative tasks, but not both. Only the contrastive image-text (i.e. CLIP-style) models can perform a wide range of tasks but rely on retrieval for the generative tasks, thus showing poor performance on these tasks. Overall, our model ChEX shows excellent performance on a wide range of tasks, covering both localization as well as text generation, and is capable of replacing specialized models without major performance drops on most tasks. Detailed results, ablation studies, and further analysis can be found in the Appendices 0.A, 0.B and 0.C.
@@
5.1.1 Localization and Region Classification Tasks
In sentence grounding (SG), ChEX demonstrates similar performance to the only generative model OmniFM-DR[82] and the best SupVG model TransVG[9] (with CheXzero[70] backbone), where TransVG shows an advantage on the mIoU metric. Notably, TransVG was trained explicitly on this task. ChEX outperforms TransVG without the CheXzero backbone by mAP and is within 1-std on mIoU. The superiority of ChEX over contrastive models underscores the importance of bounding box supervision for achieving accurate localization.
ChEX is competitive on 2 of 3 pathology detection (OD) tasks, namely NIH8 and MS-CXR. On VinDr-CXR, where bounding box supervision is available, the SupOD model Faster R-CNN [63], trained on this dataset, performs best. o-DETR[91] performs slightly worse but still outperforms ChEX, while other DETR-style models are on-par or inferior to ChEX. ChEX doubles the performance of the best WSupOD model Loc-ADPD[51] (with BioVIL backbone). On NIH8, all SupOD models perform very poor (ChEX almost doubles their best performance), while the WSupOD model Loc-ADPD is competitive with ChEX. On MS-CXR, the best SupOD baseline o-DETR (with BioVIL backbone), trained on VinDr-CXR, and the best WSupOD model Loc-ADPD are both within 1-std of ChEX. Zero-shot contrastive models perform poorly, as they rely on thresholding on noisy similarity maps.
On region classification (RC) tasks, ChEX clearly outperforms all baselines. The best baseline on MS-CXR is the Faster R-CNN with BioVIL backbone, trained on VinDr-CXR, and is outperformed by , while on CIG the best baseline is the contrastive model BioVIL, which is outperformed by .
5.1.2 Generative Tasks
Region explanation (RE) can only be performed by a single generative model, namely RGRG[69]. Other report generation models can only predict the full report but cannot provide descriptions for specified regions. On MS-CXR, our model ChEX performs similar to RGRG on Mic-F1-14 and METEOR but improves by on Mac-F1-14. On CIG, ChEX outperforms RGRG by on Mic-F1-14, by on Mac-F1-14, and by on METEOR. This result is particularly noteworthy considering RGRG’s explicit training for this task and highlights the importance of the contrastive losses in ChEX. Sentence-retrieval with contrastive baselines performs poorly on these tasks.
For full report generation (RG) on MIMIC-CXR, ChEX sets a new state-of-the-art on the Ex-F1-14, Mic-F1-5+, and Mac-F1-5+ metrics. Considering the commonly used metric Mic-F1-14, the SOTA model MAIRA-1[26] – having 7B parameters – outperforms ChEX – having only 1B parameters – by only . More notably, ChEX outperforms the more than ten times larger 12B parameter Med-PaLM M by and is only outperformed by by the 80 times larger 84B parameter version. Limitations of ChEX can be observed on the Mac-F1-14 metric, where ChEX is outperformed by the SOTA model Med-PaLM M 84B by but also by smaller models like Prompt-MRG[29] (by ) and RaDialog-RG[56] (by ). ChEX, however, improves over RGRG by , indicating that this is a limitation of training based on anatomical regions and is caused by the poor performance of some rarer pathologies (we refer to Appendix 0.C). On the Ex-F1-14 metric, ChEX outperforms the best baseline Prompt-MRG by . Most larger models do however not report this metric. On the Mic-F1-5+ and Mac-F1-5+ metrics, ChEX outperforms the large SOTA model MAIRA-1 by and , respectively. In language-based metrics such as METEOR, ChEX shows relatively low performance ( lower than the best and lower than the second-best model) due to not being explicitly trained for full report generation but generating reports based on a set of queries. While this approach offers advantages (see Sec. 5.4), it may result in reports differing from reference reports. However, it’s important to note that language-based metrics offer limited insights into the factual correctness of generated reports [49]. Overall, ChEX showcases strong report generation performance, achieving results that are close or even better than SOTA models of up to 80 times the size of ChEX.
@@ 5.4 Interpretable and Customizable Report Generation
ChEX supports automatic full report generation using a pre-defined set of textual prompts based on which the model localizes and describes relevant regions. Unlike typical report generation models, the reliance on prompt sets for report generation offers high flexibility as the prompt sets can be customized without the need for re-training. We study (cf. Appendix 0.A) the effect of using only pathologies, anatomical regions, or both as prompt sets. The choice of the prompt set enables balancing precision and recall of the model. While all studied prompt sets lead to results competitive with the baselines, with Mic-F1-14 ranging from to , using both prompt sets leads to optimal performance.
@@ 6.0.3 Limitations
While our training approach enables effective utilization of the available datasets, it also leads to some limitations. The used datasets do not provide question-answer pairs, so textual queries are either based on pre-defined prompts or report sentences, and answers are always based on report sentences. This limits the types of supported textual queries, mainly regional or pathology hints, or both. At the same time, answers are plain descriptions of the queried regions or pathologies, while specific answers to more complicated questions (e.g. comparing regions) are not supported. Also, using report sentences as answers can lead to the hallucination of comparisons with previous images, although only a single image is used, a phenomenon common to report generation models [26, 2, 61]. Future work may use instruction tuning to tackle these issues.
@@ 0.A.1.1 Token Types
Overall, this shows the importance of using all types of tokens. Having both pathology and anatomy tokens is especially relevant for good SG and OD performance, i.e. for accurate localization. Anatomy tokens are especially relevant for RC and RE tasks, due to the available of pathology labels and sentences associated with bounding boxes. Sentences tokens are the least relevant tokens and can degrade the performance on SG, i.e. they can lead to less accurate localization. However, they are relevant to achieve the best possible RE and RG performance, i.e. for optimal text generation, especially on the MS-CXR dataset and can also improve the OD performance on some datasets.
@@ 0.A.1.2 Target Types
Overall, we identified the huge importance of bounding box targets for all tasks. Pathology labels are relevant mainly for OD tasks. Again, sentence supervision is relevant for sentence generation while being less relevant for other tasks. Contrastive sentence improves some OD tasks as well as MS-CXR-based RC and RE tasks, i.e. it helps with the understanding of pathology regions.
0.A.2 Prompt Sets for Report Generation

### v2

@@ Abstract
Report generation models offer fine-grained textual interpretations of medical images like chest X-rays, yet they often lack interactivity (i.e. the ability to steer the generation process through user queries) and localized interpretability (i.e. visually grounding their predictions), which we deem essential for future adoption in clinical practice. While there have been efforts to tackle these issues, they are either limited in their interactivity by not supporting textual queries or fail to also offer localized interpretability. Therefore, we propose a novel multitask architecture and training paradigm integrating textual prompts and bounding boxes for diverse aspects like anatomical regions and pathologies. We call this approach the Chest X-Ray Explainer (ChEX). Evaluations across a heterogeneous set of 9 chest X-ray tasks, including localized image interpretation and report generation, showcase its competitiveness with SOTA models while additional analysis demonstrates ChEX’s interactive capabilities. Code: https://github.com/philip-mueller/chex.
@@
Radiology Report Generation Vision-Language Modeling1 Introduction
@@ 2.0.2 Prompt-based Localization
DETR [6] pioneered the use of token vectors for object detection, sparking a series of subsequent models following this approach [95, 94, 41, 47, 88]. However, these models typically employ tokens that are not input-dependent, often relying on learned or position-based tokens [41, 88, 47]. Visual grounding models [9, 8, 11, 37] predict bounding boxes for given textual phrases, relating them to image regions, and have been applied in medical imaging [7, 27]. GLIP [35, 89] unifies object detection and visual grounding using textual prompts, with applications emerging in the medical domain [17, 82]. Open-vocabulary object detection extends object detection models to unseen classes using textual prompts [87, 92, 93, 81, 16, 86, 42, 63, 46]. Similarly, prompt-based segmentation techniques have been explored [33] with applications in the medical domain [25, 45].
@@ 3 ChEX: A Localized and Interactive Chest X-ray Description Model
Our model ChEX supports a wide range of tasks, spanning from localization to text generation, while considering different aspects like anatomical regions and pathologies. To facilitate efficient training and zero-shot inference across all tasks, ChEX employs a simple yet versatile architecture, outlined in Figure 2. First, an image encoder is used to extract patch tokens of the given chest X-ray, while each textual prompt is encoded using the prompt encoder, a frozen text encoder. Next, the prompt detector, a DETR[6]-style object detector, localizes the prompts within the image, predicting a set of bounding boxes for each prompt along with a single Region-of-interest (ROI) token vector per prompt. Finally, the sentence generator is conditioned on each ROI token independently as well as on all patches to predict a concise description of each ROI. Further details are given in the following Secs. 3.1, 3.2 and 3.3. For reproducibility, we provide comprehensive implementation details in the supp. material.
@@ 3.2 Multitask Training
We train ChEX in a multitask setting, where each sample provides one or more types of targets, including bounding boxes for pathologies or anatomical structures, pathology classification labels (per sample or per region), and report sentences (per sample or per region). To enable training with such a wide range of targets, we use three types of prompt tokens:
1.
[…]
2.
[…]
3.
Sentence tokens: We encode each individual sentence of the radiology report provided with the current sample.
For each sample, we only use the token types for which there are targets available. Using the encoded prompt tokens, the prompt detector predicts bounding boxes and ROI tokens for each of them, before the sentence generator (conditioned on the ROI tokens) predicts the target sentences.
ChEX does not only support textual queries, i.e. prompts, but also bounding box queries. Therefore, in some (randomly selected) batches, we train ChEX with query bounding boxes instead of prompt tokens. In such cases, we compute the ROI tokens based on the target bounding boxes directly using Gaussian ROI pooling, skipping the textual tokens and box prediction process.
3.2.1 Loss Functions
We use the following loss functions to train our model:
1.
Bounding boxes: We use a modified DETR loss [6], applying Hungarian matching independently per pathology or anatomy token until all predicted boxes are matched. We retain the L1 and gIoU losses, but omit the cross-entropy loss. Box scores are instead trained only for pathologies, using the focal loss [39] with positive targets for boxes matched in the first iteration.
2.
[…]
3.
Report sentences: We apply autoregressive language modeling to predict region sentences (anatomy token) or to reconstruct the sentence (sentence token). For sentence tokens, we additionally use contrastive learning between ROI tokens and their corresponding textual sentence features. We also apply the CLIP loss [59] on average-pooled image patch and sentence features.
@@ 4.2 Benchmark Tasks
We evaluate the zero-shot performance of our model ChEX across the 9 chest X-ray tasks shown in Tab. 1. For details on the evaluation and dataset preparation, we refer to the supp. material. While no task-specific fine-tuning is applied, post-processing (e.g. box suppression and scaling of boxes) for ChEX and all baselines is adjusted to consider differences in annotation practices of datasets.
@@ 5 Results
We compare ChEX with SOTA models (Sec. 5.1). Additionally, we study how ChEX reacts to (interactive) user queries (Secs. 5.2 and 5.3), show its interpretable and customizable report generation capabilities (Sec. 5.4), and study technical design choices (Sec. 5.5). ChEX is competitive with SOTA models while providing a high degree of interpretability and interactivity, therefore offering a promising path towards clinical application.
@@ 5.1 Comparison with SOTA
Fig. 3 and Tab. 2 provide an overview of the performance of our model ChEX compared to the best baselines, including specialized SOTA and common multitask models. On 8 of the 9 tasks, ChEX is competitive (within 1-std) or better than the best baseline on at least one metric. Only on pathology detection (OD) on VinDR-CXR is it outperformed by a specialized supervised object detection model. Note that none of the baselines is capable of performing all the tasks, as most of them are either focused solely on localization or generative tasks, but not both. Only the contrastive image-text (i.e. CLIP-style) models can perform a wide range of tasks but rely on retrieval for the generative tasks, thus showing poor performance on these tasks. Overall, our model ChEX shows excellent performance on a wide range of tasks, covering both localization as well as text generation, and is capable of replacing specialized models without major performance drops on most tasks.
@@
5.1.1 Localization and Region Classification
In sentence grounding (SG), ChEX performs similarly to the generative model OmniFM-DR and the SupVG model TransVG (with TransVG showing an advantage in mIoU), despite TransVG being trained explicitly on this task. In pathology detection (OD), ChEX is competitive on 2 out of 3 tasks. On VinDr-CXR, the SupOD model Faster R-CNN performs best, while on NIH8, ChEX almost doubles the performance of the best SupOD model. On MS-CXR, ChEX is within 1-std of the best SupOD and WSupOD models. Zero-shot contrastive models perform poorly, relying on thresholding of noisy similarity maps. On region classification (RC) tasks, ChEX outperforms all baselines, with improvements of 8% on MS-CXR and 5% on CIG.
5.1.2 Text Generation
On Region explanation (RE) tasks, ChEX performs similar or better than RGRG. On MS-CXR, ChEX improves by 25% on Mac-F1-14. On CIG, ChEX improves by 18% on Mic-F1-14, 40% on Mac-F1-14, and 29% on METEOR, although RGRG was explicitly trained on this task. Other report generation models cannot provide region-level descriptions while sentence-retrieval with contrastive baselines performs poorly on these tasks.
For full report generation (RG) on MIMIC-CXR, ChEX sets a new state-of-the-art on the metrics Ex-F1-14 (), Mic-F1-5+ (), and Mac-F1-5+ (). On the commonly used Mic-F1-14 metric, ChEX outperforms the 10 times larger model Med-PaLM M 12B and is only slightly outperformed by the 7 times larger SOTA model MAIRA-1. While ChEX shows limitations on Mac-F1-14, it still outperforms RGRG. However, ChEX performs relatively low on language-based metrics like METEOR due to its query-based report generation approach. Overall, ChEX demonstrates strong report generation performance, achieving results that are close or even better than SOTA models of up to 80 times the size of ChEX.
@@ 5.4 Interpretable and Customizable Report Generation
ChEX supports automatic full report generation using a pre-defined set of textual prompts based on which the model localizes and describes relevant regions. Unlike typical report generation models, the reliance on prompt sets for report generation offers high flexibility as the prompt sets can be customized without the need for re-training. We study (cf. supp. material) the effect of using only pathologies, anatomical regions, or both as prompt sets. The choice of the prompt set enables balancing precision and recall of the model. While all studied prompt sets lead to results competitive with the baselines, with Mic-F1-14 ranging from to , using both prompt sets leads to optimal performance.
@@
5.5 Technical Insights
We additionally conducted ablation studies on several technical design decisions of our model and training paradigm. We summarize the key findings in the following. For detailed results we refer to the supp. material.
5.5.1 The mixture of tokens for pathologies, anatomy, and report sentences enables multitask capabilities.
We found that using all these three token types during training achieves the best multitask performance. Having both pathology and anatomy tokens is especially relevant for accurate localization (SG and OD tasks). Anatomy tokens are especially relevant for RC and RE tasks, likely due to the availability of pathology labels and sentences associated with bounding boxes. While sentence tokens can slightly harm localization quality, they are relevant to achieve the best possible text generation performance.
5.5.2 Localization targets are highly beneficial for all tasks, even for full report generation.
We identified the importance of bounding box targets for all tasks. Pathology class labels are relevant mainly for OD tasks. Contrastive sentence supervision improves some OD tasks as well as MS-CXR-based RC and RE tasks, i.e. it helps with the understanding of pathology regions, while generative text supervision is not significantly beneficial for non-generative tasks.
5.5.3 Technical differences to baselines
A key distinction between ChEX and most generative baselines lies in its aspect-level generation approach, where sentences are generated individually for specific findings and regions of the image. In contrast, models like MAIRA-1[26], Med-PaLM M [72], or OmniFM-DR[83] generate the full report in a single shot. This approach is critical to ChEX’s strong performance, as demonstrated by the good results of RGRG[70], which employs a similar region-level strategy. Compared to RGRG, ChEX introduces two key innovations: (i) pathology and sentence tokens, which extend RGRG’s solely anatomy-based approach; and (ii) contrastive alignment of region features with their sentences and pathology prompts. Our ablation studies confirm that these differences are essential to ChEX’s superior performance over RGRG.
@@ 6.0.3 Limitations
Our training approach effectively utilizes the available datasets and eliminates the need for intricate data engineering. However, this comes at the cost of a more complex training process. Furthermore, the used datasets do not provide question-answer pairs, so textual queries are either based on pre-defined prompts or report sentences, and answers are always based on report sentences. This limits the types of supported textual queries, mainly regional or pathology hints, or both. At the same time, answers are plain descriptions of the queried regions or pathologies, while specific answers to more complicated questions (e.g. comparing regions) are not supported. Also, using report sentences as answers can lead to the hallucination of comparisons with previous images, although only a single image is used, a phenomenon common to report generation models [26, 2, 62]. Future work may use instruction tuning to tackle these issues.
Additionally, as we move towards clinical application, further evaluation from a radiologist’s perspective is essential, and future work should include systematic studies on user experience to ensure seamless integration into clinical workflows.
@@ 0.A.1.2 Target Types
0.A.2 Sentence Generator
In Tab. 4, we provide an ablation study on design choices of the sentence generator. More precisely, we studied (i) the relevance of the post decoder, a component of the sentence decoder; (ii) variations of the query-key-value (qky) features used in P-tuning (linear projection instead of MLP, shared projections for all decoder layers, shorter prefix length); and (iii) the benefits of decoder language model pre-training on MIMIC-CXR. Overall, we found that most design decisions, except the post decoder, have little impact.
Sentence Generator | MS-CXR [Mic-F1-14] | CIG [Mic-F1-14] | MIMIC-CXR [Mic-F1-14] |
ChEX | 49.97 | 53.34 | 52.32 |
no post decoder | 45.86 | 43.90 | 46.07 |
linear qkv proj. | 48.64 | 52.80 | 52.41 |
layer-shared qkv | 48.25 | 53.38 | 52.46 |
prefix length 1 | 49.65 | 53.14 | 52.25 |
no gen. pretraining | 48.56 | 52.89 | 52.46 |
We argue that the post decoder is so beneficial, because it enables further processing of ROI features even if only bounding boxes where given as queries (where huge parts of the detector are skipped). The choice of the projections during P-tuning has little influence, such that we opted for the default configuration (MLP projections without sharing). While the prefix length has minimal influence, a slightly longer prefix length of 5 (as used in CHeX), provide some minor improvements across all generation tasks. Skipping pre-training of the decoder on MIMIC-CXR and directly using the decoder language model pre-trained only on PubMed, leads to small performance degradation on some tasks. However, the main benefit of this pre-training is a faster convergence in the final end-2-end training.
0.A.3 Query Type for Region Classification and Explanation
Chest ImaGenome (CIG) | ||
Region Classification (RC) | Region Explanation (RE) | |
Query type | [wAUROC] | [Mic-F1-14] |
Boxes (default task) | 70.460.36 | 53.340.43 |
Textual prompt | 70.760.37 | 54.700.41 |
0.A.4 Prompt Sets for Report Generation